# Conceptualization, Methodology, Software, Writing

### v1

@@ 1 Introduction
Nowadays, automatic visual recognition systems have become increasingly popular as powerful support tools for a wide range of vision-related applications, e.g., object detection and tracking [4, 9], image analysis and classification [42], or scene segmentation [37] and understanding [10]. Such systems have been developed to mimic the remarkable human brain ability to perfectly correlate low-level visual features (e.g., edges, color, or shapes) and semantic-level information for perceptual identification and recognition tasks [1, 2]. In the past decade, deep learning algorithms have been extensively employed to model how the human brain processes visual information since artificial neural networks are inspired by the structure and function of biological neurons; indeed, they can analyze visual data learning to recognize patterns and features in images [27]. Precisely, the understanding of visual content has been revolutionized by Convolutional Neural Networks (CNNs), primarily because of their ability to analyze an image split into smaller patches, extract multi-scale localized features, and synthesize them to generate highly expressive representations [21, 19]. Despite this, due to its limited receptive field, the convolutional operation of standard CNNs cannot recognize long-range dependencies, such as arbitrary relationships among distant image regions. To address this limitation and facilitate the handling of long-range dependencies in visual tasks, several architectures based on transformers have emerged [7, 33]. Notably, the traditional transformer neural network initially gained popularity in the field of natural language processing [34]. This popularity stemmed from its self-attention mechanisms, which enable the modeling of extensive dependencies between tokens within a sequence of data, overcoming the limitation of CNNs. Additionally, in contrast to CNNs, specifically engineered for visual tasks with pre-existing knowledge and incorporating inductive biases such as locality and translation equivariance [8], ViT adopts an approach where an image is treated as a sequence of patches. This different architectural choice leads to a deficiency in such inductive biases, making ViT less effective at generalizing when trained on smaller datasets [7]. Nevertheless, it has been observed that the training scale can override the need for strong inductive biases, showcasing the substantial potential of transformers when trained on extensive datasets [7, 8, 33].
@@
Standard ViT models are characterized by their deep architecture, comprising multiple sequential computational blocks. Each of these blocks consists of a self-attention layer, a multi-layer perceptron (MLP), and a residual connection that bypasses the self-attention layer. This hierarchical structure necessitates multiple consecutive passes to process the input token sequence and compute scene representations. While the input token sequence traverses deeper layers of the network, tokens tend to become highly similar to each other. This phenomenon leads to a loss of feature diversity, commonly referred to as feature collapsing. The main contributing factor to this phenomenon is the global nature of the self-attention mechanism and its tendency to combine features across patches [32]. Specifically, as token processing progresses through the network’s deeper layers, the generated attention maps become increasingly global, aggregating features between patches at a global scale and losing their focus on local information [7, 8]. This shift from local awareness causes the network to forget low-level features and only emphasize global characteristics. While the ability to capture global context is a desirable trait of transformers, considering low-level cues such as shape and color, typically found in shallower network layers, can be advantageous for vision-related tasks like image classification and object detection. In addition to the globality issue, ViT architectures are also considered less robust at translation invariance compared to CNNs [16, 7]. This limitation makes ViTs less reliable in real-world scenarios where objects may appear in varying positions and sizes.
Inspired by the shortcomings observed in the ViT architecture and insights from neuroscience, which suggest that the human brain transmits visual information through a sequential hierarchical process from simple to specialized cells via the aggregation of electrical impulses [15], we present the Residual Attention Vision Transformer (ReViT) [^1] https://github.com/ADiko1997/ReViT. This innovative network architecture introduces an attention residual learning technique, enhancing the performance of visual recognition tasks by integrating low-level information from shallow to deeper network layers and contrasting the over-globalization of the attention mechanism. Specifically, given standard vision transformer blocks, attention information is passed between consecutive layers and integrated with the attention information of the current layer using a permutation invariant aggregation function. This connection facilitates the transmission and accumulation of attention-related information from shallower to deeper layers, enabling the ability to identify low-level features while preserving the original ability of transformers to extract global context. Consequently, by empowering the attention mechanism with the ability to capture such features, the resulting learned representations are expected to exhibit greater feature diversity. Additionally, since low-level features like color and shape can be useful in identifying objects in a scene regardless of their position and size, this mechanism improves the translation invariance of ViT. In order to assess the efficacy of incorporating residual attention into ViT (i.e., ReViT), we conducted a comprehensive empirical analysis. Our evaluation spanned five image classification benchmarks, namely ImageNet1K [30], CIFAR-10 [18], CIFAR-100 [18], Oxford Flowers-102 [28], and Oxford-IIIT Pet [29]. The results demonstrated significant performance enhancements when using ReViT compared to the original ViT counterpart approach in all datasets. Furthermore, due to their high image quality, we employed the Oxford Flowers-102 and Oxford-IIIT Pet datasets to highlight ReViT superiority over ViT in image classification tasks involving translation invariance, such as horizontal and vertical shifts, as well as scale invariance. These experiments provided quantitative evidence of ReViT advantages in handling these scenarios. Moreover, we quantitatively analyze the global spread of the attention map in ViT and ReViT models employing the non-locality metric [8] and show that our approach can preserve more local information than ViT. Apart from our quantitative evaluations, we conduct a qualitative analysis by comparing the feature maps learned by ViT and ReViT models employing the GradCAM [31] algorithm to visually demonstrate how ReViT excels in integrating low-level features into its learned representations, as depicted in Fig. 1. Finally, we established that the proposed residual learning strategy can seamlessly integrate into multiscale transformer architectures like the Multiscale Vision Transformer v2 (MViTv2) [20] and the Shifted-window transformer (Swin) [24], verifying its effectiveness in various tasks, including image classification on ImageNet1K, object detection, and instance segmentation on COCO2017 [23], obtaining improved performances.
@@
    To the best of our knowledge, a novel ViT-based architecture has been introduced that uses residual attention modules capable of incorporating important low-level visual characteristics into the learned representation while maintaining the ability to extract global context, thereby enhancing feature diversity throughout deeper network layers.
@@
    We conduct an extensive experimental phase to showcase the effectiveness of the proposed model in image classification on benchmarks like ImageNet1K, CIFAR-10, CIFAR-100, Oxford Flowers-102, and Oxford-IIIT Pet.
@@
    Empirically demonstrate that residual attention enhances the robustness of ViT against translation invariance through a comprehensive evaluation in the context of image classification tasks on the Oxford Flowers-102 and Oxford-IIIT Pet datasets.
  4. 4.
    Thoroughly analyze the global spread of attention maps, with and without residual attention, using the non-locality metric.
  5. 5.
    We qualitatively analyze the learned feature maps of ViT and ReViT models using the GradCAM algorithm to examine the attention mechanism’s ability to integrate low-level features into acquired representations in the presence of residual attention.
  6. 6.
    We evaluate our module’s seamless integration into multi-scale architectures using MViTv2 and Swin with residual attention on ImageNet1K for image classification and COCO2017 for object detection and instance segmentation.
The rest of this paper is organized as follows. Section 2 introduces relevant work that inspired this study. Section 3 comprehensively describes important background information and the proposed method. Section 4 describes the experiments performed to validate the proposed approach quantitatively and qualitatively. Section 5 reports an ablation study on the residual attention. Finally, Section 6 draws some conclusions from this study.
@@ 2.1 Vision Transformers
In visual recognition tasks, standard CNN-based models are currently the most powerful algorithms for image analysis [14, 35]; even so, these networks are designed to capture short-range dependencies using small convolutional filters that only look at small image patches at a time. However, due to its self-attention mechanisms, which have been very successful for natural language processing, the transformer architecture was recently applied to process images, capturing long-range dependencies between image patches, thus allowing the handling of complex visual relationships better. Therefore, the ViT [7] neural model was introduced, splitting the image into a grid of patches fed through a stack of transformer blocks, which use self-attention to encode the global relationships between the different patches. To enhance ViT capabilities, Touvron et al. [33] proposed the Data-efficient Image Transformers (DeIT) using a knowledge distillation technique that transfers the knowledge learned by a larger teacher model to a smaller student model (the DeiT model) during training. This helps to improve the generalization performance of ViT. Despite the promising results, extracting low-level image features becomes a significant challenge. To address such an issue, Han et al. [11] introduced a new kind of neural architecture that divides image patches into sub-patches and later employs an extra transformer block to analyze their relationship. For the same reason, Liu et al. [24] presented the Swin transformer, a multiscale architecture operating in local windows with a shifted window policy to extract spatial cross-window relationships between patches. Moreover, in literature, several works proposed enhancing local feature representations by aggregation strategy [39] and integrating local and global attention layers [5, 22]. Aiming to extract features at different levels of granularity and resolution, Li et al. [20] proposed the MViTv2 to capture local and global image features using a hierarchical architecture with multiple scales, which can improve the model ability to define and recognize objects within a scene in different sizes and orientations. Differently, to improve the performance of ViT-based models for large-scale image classification, Wang et al. [36] introduced the kNN search inside the attention mechanism to reduce the impact of noisy or irrelevant features by exploiting the locality of patches to extract meaningful information, thus resulting in overall better feature representation. Finally, Yu et al. [38] proposed the Bilateral lOcal Attention vision Transformer (BOAT) based on an attention mechanism that integrates the common image-space local attention with feature-space local attention. The latter can compute attention among relevant image patches even if they are not close to each other within the image plane. Therefore, this type of attention is a natural compensation to image-space local attention, which may miss capturing meaningful relationships between existing patches across different local windows. However, kNN-based and BOAT models require significant computational resources, limiting their practical applicability.
Due to their success in increasing CNNs capabilities, ViT and all its variants include residual connections. However, their usual application does not completely solve the feature collapsing of ViT in deeper layers. Therefore, addressing such an issue, we propose residual attention learning for vision transformers, which propagates attention among adjacent transformer blocks while preserving the computational load, improving feature diversity and performances in tasks like image classification, object detection, and instance segmentation.
2.2 Feature Diversity
@@
Traditional downstream ViT models are characterized by deep architectures, resulting in reduced feature diversity as the depth increases. This limitation hampers their ability to represent information effectively. While variants of ViT that have similarities to CNNs, e.g., MViTv2 or Swin, are less affected by this issue, it should be considered that they are primarily designed for image-specific tasks, deviating from the original ViT goal of achieving generalization across a wide range of tasks and modalities. Consequently, addressing this challenge is essential for enhancing the representational capabilities of standard ViT models and promoting consistency in hardware and software designs across various tasks and modalities. To tackle this problem, Tang et al. [32] found inspiration from residual connections, which aids in maintaining a degree of feature diversity. They propose an augmented shortcut scheme that enhances the flow of information from low-level layers to deeper layers through multiple parallel alternative paths. These shortcuts are essentially parametrized linear projection sequences that bypass the attention mechanism within the transformer block. This augmentation significantly boosts feature diversity as the information travels through the deeper layers of ViT. In contrast, d’Ascoli et al. [8] introduced a novel attention layer called the gated positional attention layer. When equipped with a soft inductive bias to mimic the behavior of CNNs, this layer improves the feature diversity of ViT models. However, these methods either require the addition of extra parameters or alterations to the standard ViT workflow by changing the composition of the transformer block. For these reasons, the proposed method offers an elegant solution. It increases feature diversity in ViT architectures by forwarding attention among transformer layers with minimal additional parameters and seamlessly integrating into the original ViT workflow. Additionally, contrary to [32], the proposed method addresses the issue of feature diversity by contrasting attention over-globalization, which is the main catalyst in feature collapsing rather than propagating past features.
@@ 3.1 Vision Transformer Layer
The standard ViT adopts the typical architecture as proposed in [34] with MLP and attention modules alternatively stacked throughout the network. Among different types of attention mechanisms, scaled dot-product self-attention is the most used one. It enables the model to extract complex relationships among the elements in the input data sequence and to dynamically assign varying significance degrees to each element according to the learned relationships [34]. Formally, given a set of image patch features $X=\{x_{1},x_{2},x_{i},...,x_{N}\}$, with $X\in R^{N\times C}$, where $x_{i}$ is the vector representation of the $i$-th image patch, $N$ is the total number of patches, and $C$ is the channel size of the image, a learnable linear projection layer is used to generate query $Q$, key $K$, and value $V$ vectors, such that $Q,K,V\in R^{N\times d}$ with $d$ being the projection dimensionality. Afterward, for a given attention layer $l$, the dot product between queries and keys is computed and scaled by the square root of $d$ to produce raw attention scores vector $S_{l}$, as shown in Fig. 2(a). Therefore, the softmax function is applied to such scores, computing the attention weights matrix $A_{l}$, as follows:
@@
Finally, the values in $V_{l}$ are scaled using the attention weights through a weighted multiplication to compute the scaled dot-product attention, defined as:
$\tilde{X}_{l}=A_{l}V_{l},$ (3)
where $\tilde{X}_{l}$ represents the output features of the scaled dot-product attention at layer $l$.
However, due to the global nature of the self-attention mechanism, processing a considerable number of patches or a big $d$ requires a high computational cost. Therefore, Ashish et al. [34] proposed the Multi-Head Self-Attention mechanism (MHSA) to retain the computational requirements when a significant workload is required. This approach uses $H$ different attention heads computing self-attention in different subspaces with dimension $d/H$ derived from the Q, K, and V vectors rather than considering their full dimensionality $d$. In this manner, the calculations are parallelized, and the model can jointly attend to information from different representations of subspaces at different positions. Finally, the output of each head is concatenated together to obtain the final attention output. Formally, the MHSA operation is represented as:
$MHSA(X_{l})=Concat([A_{l,h}V_{l,h}]^{H}_{h=1}),$ (4)
where $h$ represents the head index, $A_{l,h}$ and $V_{l,h}$ are the attention weights matrix and the value vectors of head $h$ and layer $l$ respectively, while $Concat(\cdot)$ denotes the concatenation of the $H$ produced feature maps along the head dimension. Consequently, $\tilde{X}$ can now be expressed as follows:
$\tilde{X}_{l}=Concat([\tilde{X}_{h,l}]^{H}_{h=1}),$ (5)
where $\tilde{X}_{h,l}$ represents the attention scaled values at layer $l$ and head $h$.
The MLP module, on the other hand, extracts features independently from each patch and is usually constructed by stacking two linear layers with parameters $W_{a}$, $W_{b}$ respectively, and a non-linear activation function $\sigma$ in between them. This module also represents the last processing operation of the transformer block, which takes in input $\tilde{X}_{l}$ and produces the output features of layer $l$ denoted as $X_{l+1}$ since are going to be used as the input features for the $l+1$ layer. Formally, it can be represented as:

### v2

@@ 1 Introduction
Nowadays, automatic visual recognition systems have become increasingly popular as powerful support tools for a wide range of vision-related applications, e.g., object detection and tracking [2, 8], image analysis and classification [46], or scene segmentation [41] and understanding [10]. Such advancements were initially fueled by the prowess of Convolutional Neural Networks (CNNs) in extracting multi-scale localized features [23, 20]. However, CNNs are bounded by limited receptive fields, hindering their ability to model arbitrary relationships among distant image regions. To address these limitations, several works started exploring vision transformer (ViT) architectures [6, 36] built upon the self-attention mechanisms, which unlocks the modeling of extensive dependencies between tokens within a data sequence [37]. Adapting with such a mechanism, in contrast to CNNs, engineered for visual inputs with built-in inductive biases such as locality and translation equivariance [7], ViT treats images as a sequence of local patches. However, despite using such an unconventional approach, ViT has propelled computer vision to new heights, showcasing the potential of a unified architecture for diverse tasks and modalities [9].
@@
While ViTs have revolutionized computer vision, their reliance on vast training datasets (hundreds of millions to billions of samples [6]) limits their applicability to smaller datasets impeding their development for domain-specific applications [21]. Such reliance has effectively rendered datasets like ImageNet1K [32], formerly considered a rigorous benchmark for training vision models, a legacy standard primarily used for assessing classification performance. However, even though modern computing power allows for large-scale training, the ability to converge on smaller datasets remains crucial. This is especially important for enabling real-world applications dealing with specialized domains like medical imaging [47], aerial imaging [19], etc., where data collection is expensive and limited. Moreover, data in these domains often diverges significantly from the data used to pre-train large transformer models, hindering their effectiveness [21]. These limitations introduce important challenges to ViT architectures, emphasizing the need for modern solutions that can also unlock its ability to converge in smaller-scale datasets.
According to the literature, the primary rationale for such limitation is twofold: 1) the absence of inherent inductive bias [7, 6, 26], and 2) feature collapsing [35]. The former arises from the self-attention mechanism, which is explicitly designed for images and, therefore, lacks inherent assumptions about the spatial structure of the data. However, despite hindering the performance of transformers on restricted data, this can also be considered a treat since it is crucial in enabling the flexibility to generalize across modalities. On the other hand, the second and most critical rationale, feature collapsing, refers to the degradation of information within transformer architectures as features progress through the transformer network. Specifically, transformers are typically composed of stacked computational blocks, including multi-head self-attention layers (MHSA), multi-layer perceptron (MLP), layer normalization operators (LN), and residual connections. As the input features flow forward through these blocks, they become indistinguishable due to inter-token feature aggregation within the self-attention mechanism [35]. This causes feature collapsing, resulting in overly-global representations that encode only global information, omitting crucial local details. Such a phenomenon hinders the network’s ability to extract robust representations necessary for vision tasks like object classification or detection [36, 35].
In their attempts to improve the vision transformer architecture, many prior works have proposed interesting solutions by implicitly or explicitly targeting the above limitations. For instance, in [36], authors propose using heavy data augmentation and knowledge distillation from CNNs to improve learning efficiency. Following, several other works try to inspire from the effectiveness of CNNs and morph the transformer architecture by inducing CNN-like initialization bias [7], windowed processing [26], and multi-scale hierarchical transformer architectures [22, 26] able to mimic the convolution operation using the attention mechanism. However, while these approaches might be effective, they diverge from the natural ability of transformers to generalize across modalities and enable the unification of software and hardware design. In such regards, in [35], authors propose to improve the efficiency of transformers targeting features collapsing through augmented residual connections between features of consecutive attention blocks. This is the first work addressing the core issue of transformers related to learning efficiency. Nevertheless, the proposed approach introduces a non-negligible computational burden as it requires the addition of several MLP layers parallel to each transformer block, making it less suitable for real-world applications. Moreover, such an attempt fails to address the feature collapsing at its core: self-attention over-globalization.
In pursuit of addressing these transformer limitations and the shortcomings of previous approaches, this work presents the Residual attention Vision Transformer (ReViT) [^1] https://github.com/ADiko1997/ReViT architecture. Inspired by the human brain’s hierarchical processing of visual information, where simple cells aggregate electrical impulses to transmit signals to specialized cells [16], ReViT tackles the challenge of feature collapsing observed in conventional ViTs. It does so by introducing a novel attention residual learning technique contrasting the over-globalization of self-attention, a phenomenon that fuels feature collapsing as highlighted in [35, 6, 7]. Specifically, given standard vision transformer blocks, attention information is passed between consecutive layers and integrated with the attention information of the current layer in the form of a residual connection using a permutation invariant aggregation function. This connection facilitates the transmission and accumulation of attention-related information from shallower to deeper layers, enabling the ability to identify low-level features while preserving the original ability of transformers to extract global context. Consequently, by empowering the attention mechanism with the ability to capture such features, the resulting learned representations are expected to exhibit greater feature diversity. Additionally, due to its simple nature, ReViT does not morph the core transformer architecture preserving its innate capabilities and computational load.
In order to assess the efficacy of incorporating residual attention into ViT (i.e., ReViT), we conducted a comprehensive empirical analysis. Our evaluation spanned five standard image classification benchmarks, namely ImageNet1K [32], CIFAR-10 [18], CIFAR-100 [18], Oxford Flowers-102 [30], and Oxford-IIIT Pet [31]. These datasets represent a range of scales and complexities, including medium-to-small scale datasets that better reflect real-world scenarios with limited data availability. Our findings demonstrate consistent performance improvements with ReViT compared to the standard ViT across all benchmarks, even under challenging conditions of image alterations that test the model’s robustness. Moreover, quantitative analysis of the attention maps, using the non-locality metric [7], reveals that ReViT effectively preserves more local information than standard ViTs, showcasing its effectiveness in preserving feature diversity. Further qualitative analysis visually illustrates ReViT’s superior ability to integrate low-level features into its learned representations, as depicted in Fig. 1. Finally, we established that the proposed residual learning strategy can seamlessly integrate into multiscale transformer architectures like the Multiscale Vision Transformer v2 (MViTv2) [22] and the Shifted-window transformer (Swin) [26], verifying its effectiveness in various tasks, including image classification on ImageNet1K, object detection, and instance segmentation on COCO2017 [25], obtaining improved performances.
@@
    This work presents ReViT, a novel Vision Transformer architecture designed to tackle the limitation of feature collapsing. ReViT leverages the concept of residual attention, which, to the best of our knowledge, has not been previously explored.
@@
    Demonstrate empirically and theoretically that residual attention effectively contrasts feature collapsing on ViT architectures.
@@
    Extensive experimental results in several benchmarks assess the residual attention mechanism in enhancing performance across visual tasks (image classification, semantic segmentation, and object detection) on single and multi-scale architectures, demonstrating its effectiveness and versatility.
The rest of this paper is organized as follows. Section 2 introduces relevant work that inspired this study. Section 3 comprehensively describes important background information and the proposed method. Section 4 describes the experiments performed to validate the proposed approach quantitatively and qualitatively. Section 5 reports an ablation study on the residual attention. Finally, Section 6 draws the final conclusions from this study.
@@ 2.1 Vision Transformers
In visual recognition tasks, standard CNN-based models are currently the most powerful algorithms for image analysis [15, 38]; even so, these networks are designed to capture short-range dependencies using small convolution filters that only look at small image patches at a time. However, due to its self-attention mechanisms, the transformer architecture was recently applied to process images, capturing long-range dependencies between image patches, thus enabling better modeling of complex visual relationships. Therefore, the ViT [6] neural model was introduced, splitting the image into a grid of local patches fed through a stack of transformer blocks, which use self-attention to encode the global relationships between the different patches. To enhance ViT capabilities, Touvron et al. [36] proposed the Data-efficient Image Transformers (DeIT) using a knowledge distillation technique that transfers the knowledge learned by a larger CNN teacher model to a transformer model during training. This helps to improve ViT’s generalization performance. However, despite the promising results, extracting low-level image features becomes a significant challenge. To address such an issue, Han et al. [11] introduced a new kind of neural architecture that divides image patches into sub-patches and later employs an extra transformer block to analyze their relationship. For the same reason, Liu et al. [26] presented the Swin transformer, a multiscale architecture operating in local windows with a shifted window policy to extract spatial cross-window relationships between patches. Trying to address the shortcomings of local-window self-attention in vision transformers, such as limited receptive field and weak modeling capabilities, Qiang et al. [3] introduced the MixFormer model integrating local-window self-attention with depth-wise convolution in a parallel configuration, enhancing the model’s ability to capture both intra-window and cross-window connections. Moreover, in literature, several works proposed enhancing local feature representations by aggregation strategy [43] and integrating local and global attention layers [4, 24]. Aiming to extract features at different levels of granularity and resolution, Li et al. [22] proposed the MViTv2 to capture local and global image features using a hierarchical architecture with multiple scales, which can improve the model’s ability to define and recognize objects within a scene in different sizes and orientations. Differently, to improve the performance of ViT-based models for large-scale image classification, Wang et al. [39] introduced the kNN search inside the attention mechanism to reduce the impact of noisy or irrelevant features by exploiting the locality of patches to extract meaningful information, thus resulting in overall better feature representation. Again, Yu et al. [42] proposed the Bilateral lOcal Attention vision Transformer (BOAT) based on an attention mechanism that integrates the common image-space local attention with feature-space local attention. The latter can compute attention among relevant image patches even if they are not close to each other within the image plane. Therefore, this type of attention is a natural compensation to image-space local attention, which may miss capturing meaningful relationships between existing patches across different local windows. However, kNN-based and BOAT models require significant computational resources, limiting their practical applicability. More recently, still recognizing the limitations of existing methods that focus primarily on local or regional feature extractions, Ting et al. [40] proposed the Dual Vision Transformer (Dual-ViT) that integrates two distinct pathways: a semantic pathway for global feature compression and a pixel pathway for local detail enhancement. This design strategy optimizes the self-attention mechanism by balancing the computation between global semantic understanding and local feature refinement, reducing the computational load while enhancing the model performance. Investigating multi-scale feature modeling, Nie et al. [29] proposed ScopeViT, an efficient scale-aware vision transformer model that enhances the ViT capacity to perceive and integrate visual tokens at different scales by embedding them with varying receptive fields into distinct attention heads. This model is characterized by two lightweight attention mechanisms facilitating varied scale perception across the network and optimizing the computational cost by reducing the token count in attention processes. Again, Tang et al. [34] proposed the CATNet network employing the Multi-dimensional Convolutional Attention (MCA) module, enhancing the model’s capability to process multi-scale features, and the Dual Attention Transformer (DAT) module, improving the accuracy of pixel-level predictions. Also, this network architecture tries to maintain high accuracy while significantly reducing the computational load by integrating convolutional attention mechanisms with transformer-based approaches. This work differs from the previous ones in not seeking to make transformers resemble more CNNs but proposing a novel residual connection that helps improve feature diversity on ViTs and improve their representation capabilities.
2.2 Residual Connections and Feature Diversity
@@
Traditional downstream ViT models are characterized by deep architectures, which typically result in reduced feature diversity as the depth increases. This limitation hinders their ability to represent information effectively. Residual connections, pioneered by He et al. [14] presenting the ResNet architecture, are a well-established method to alleviate the vanishing gradient problem in deep networks and maintain feature diversity by allowing the unimpeded flow of gradients through the network. These connections create a direct pathway for earlier layers’ features to impact deeper layers, enhancing the model’s ability to learn diverse feature representations as the network goes deeper for downstream prediction. Inspired by these effects, residual connections between consecutive blocks have also been incorporated into the transformer architecture [37], and its vision homologue ViT [6]. Motivated by its impact, Tang et al. [35] proposed an augmented residual connection scheme in ViT models that enhances the flow of information between layers through multiple parallel alternative paths. These shortcuts, essentially parametrized linear projection sequences, bypass the attention mechanism within the transformer block. This augmentation boosts feature diversity on the deeper layers of ViT. In contrast, while variants of ViT that bear similarities to CNNs, such as MViTv2 or Swin, are less affected by the reduction in feature diversity, it is essential to note that these models are primarily designed for image-specific tasks, deviating from the original ViT goal of achieving generalization across a wide range of tasks and modalities. Furthermore, d’Ascoli et al. [7] introduced a novel attention layer called the gated positional attention layer in ViTs. Equipped with a soft inductive bias to mimic the behavior of CNNs, this layer improves the feature diversity of ViT models. However, while these methods contribute to increased feature diversity, they either require adding significant computational load or alterations to the standard ViT workflow by changing the composition of the transformer block. For these reasons, the proposed method offers an agile solution. It increases feature diversity in ViT architectures by forwarding attention among transformer layers with minimal additional parameters and seamlessly integrating into the original ViT workflow. Additionally, contrary to [35], the proposed method addresses the issue of feature diversity by contrasting attention over globalization, which is the main catalyst in feature collapsing rather than propagating past features.
@@ 3.1 Vision Transformer Layer
The standard ViT adopts the typical architecture as proposed in [37] with MLP and attention modules alternatively stacked throughout the network. Among different types of attention mechanisms, scaled dot-product self-attention is the most used one. It enables the model to extract complex relationships among the elements in the input data sequence and to dynamically assign varying significance degrees to each element according to the learned relationships [37]. Formally, given a set of image patch features (or tokens) $X=\{x_{1},x_{2},x_{i},...,x_{N}\}$, with $X\in R^{N\times C}$, where $x_{i}$ is the vector representation of the $i$-th image token, $N$ is the total number of tokens, and $C$ is the channel size of tokens, a learnable linear projection layer is used to generate query $Q$, key $K$, and value $V$ vectors, such that $Q,K,V\in R^{N\times d}$ with $d$ being the projection dimensionality. Once these vectors are obtained, for a given attention layer $l$, we first calculate the dot product between queries, scaling it by the square root of $d$ to prevent vanishing gradient [37] and produce raw attention scores vector $S_{l}$:
@@
Therefore, the softmax function is applied to discretize such scores in the $[0,1]$ range, resulting in the attention weights matrix $A_{l}$:
@@
These weights determine how much each token contributes to the representation of other tokens. Finally, we compute the output features $\tilde{X}_{l}$ by aggregating the value matrix $V_{l}$ using the attention weights:
$\tilde{X}_{l}=A_{l}V_{l}.$ (3)
Despite being exceptional in modeling token relationships, standard self-attention mechanisms can be computationally expensive when handling a large number of image patches or high-dimensional features. To tackle this, the MHSA approach, introduced in [37], provides an elegant solution. Instead of processing the full-dimensional $Q$, $K$, and $V$ vectors directly, MHSA splits them into $H$ smaller subspaces named heads, each with a reduced dimension of $d/H$. This enables parallel calculations across these subspaces, allowing the model to simultaneously focus on different aspects of the input information. Formally, for a given head $h$ and layer $l$, the output attention features $\tilde{X}_{l,h}$ can be obtained as follows:
$\tilde{X}_{l,h}=A_{l,h}V_{l,h},$ (4)
where $h$ represents the head index, $A_{l,h}$ and $V_{l,h}$ are the attention weights matrix and the value vectors of head $h$ and layer $l$, respectively. The outputs from each head are then combined, providing a comprehensive attention result that incorporates the insights gained from each individual subspace. Formally, according to the definition of MHSA, $\tilde{X}_{l}$ can be represented as:
$\tilde{X}_{l}=Concat([A_{l,h}V_{l,h}]^{H}_{h=1}),$ (5)
where $Concat(\cdot)$ denotes the concatenation of the $H$ produced feature maps along the head dimension to provide the comprehensive MHSA results in $\tilde{X}_{l}$. The MLP module, on the other hand, extracts features independently from each patch and is usually constructed by stacking two linear layers with parameters $W_{a}$, $W_{b}$ respectively, and a non-linear activation function $\sigma$ in between them. This module also represents the last processing step of the transformer block, which takes in input $\tilde{X}_{l}$ and produces the output features of layer $l$ denoted as $X_{l+1}$ since they are going to be used as the input features for the $l+1$ layer. Formally, it can be represented as: